# Majorization via positivity of Jack and Macdonald polynomial differences

### v1

@@ 5. Macdonald polynomial differences: conjectures and results
    For some fixed $q_{0},t_{0}\in(1,\infty)$, we have
@@
    For some fixed $q_{0},t_{0}\in(1,\infty)$, we have
@@
By the well-known duality $P_{\lambda}(x;q,t)=P_{\lambda}(x;1/q,1/t)$, one can define instead $\mathbb{F}_{\geqslant 0}\coloneqq\{\,f\in\mathbb{Q}(q,t)\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}f(q,t)\geqslant 0,\text{\ if\ }0<q,t<1\,\}$ and get an equivalent statement.
Remark 5.5.
For completeness, we record two additional equivalent conditions in Conjecture 6:
  • $(1^{\prime})$
    For all real $q,t>1$, the Macdonald difference lies in the Muirhead semiring over $\mathbb{R}_{\geqslant 0}$:
  • $(2^{\prime})$
    For all real $q,t>1$, we have
    $\frac{P_{\lambda}(x;q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x;q,t)}{P_{\mu}(\bm{1};q,t)}\in[0,\infty),\quad\forall x\in[0,\infty)^{n}.$ (5.10)
These come from specializing assertions (1) and (2) respectively, to scalars $q,t>1$. It is also clear that $(1)\implies(1^{\prime})\implies(2^{\prime})$ and $(1)\implies(2)\implies(2^{\prime})$; and $(2^{\prime})\implies(3)$. Moreover, the counterpart of $(2^{\prime})$ can also be added in Conjectures 1 and 2 (and in Conjecture 7); as also the counterpart of $(1^{\prime})$ in Conjecture 3.
@@
Each assertion in Conjecture 6 implies the next. Moreover, the implication $(5)\implies(1)$ – and hence Conjecture 6 – moreover holds for all $\lambda$, with $\mu=(1^{|\lambda|})$. It also holds for $n=2$ and $\lambda,\mu$ arbitrary.
@@
That $(1)\implies(2)$ follows from Muirhead’s inequality, and $(2)\implies(3)\implies(4)$ are obvious. The proof of $(4)\implies(5)$ is via the argument used to show Theorem 3.2: if (4) holds for all $x\in(1,\infty)^{n}$ then $\lambda$ weakly majorizes $\mu$; if it holds for $x\in(0,1)^{n}$ then $\lambda\succcurlyeq_{w}-\mu$. Together, these are equivalent to $\lambda\succcurlyeq\mu$.
Next, we assume $\mu={\bf 1}_{|\lambda|}$, with $n,\lambda$ arbitrary and show $(5)\implies(1)$. Haglund–Haiman–Loehr have shown [7] that the integral Macdonald polynomials are monomial-positive over $\mathbb{F}_{\geqslant 0}$ for $q,t\in(0,1)$, whence so are the $P_{\lambda}$. Thus by the duality in Section 5, $P_{\lambda}(x)$ is monomial-positive for $q_{0},t_{0}\in(1,\infty)$. Now the calculation in Eq. 4.2 implies (1).
Finally, let $n=2$ and $\lambda,\mu$ be arbitrary. To show $(5)\implies(1)$, it suffices via telescoping to do so for majorization-adjacent partitions $\lambda=(d-i,i)$ and $\mu=(d-i-1,i+1)$. We showed this in Section 5 for $i=0$; and we use this to compute for $i>0$:
@@
$\in(x_{1}x_{2})^{i}\cdot\sum_{\nu\mathrel{\succcurlyeq}\xi}\mathbb{F}_{\geqslant 0}\cdot(M_{\nu}(x)-M_{\xi}(x))$
$=\sum_{\nu+(i,i)\mathrel{\succcurlyeq}\xi+(i,i)}\mathbb{F}_{\geqslant 0}\cdot(M_{\nu+(i,i)}(x)-M_{\xi+(i,i)}(x)).\qed$
@@
We conclude with two other conjectures, which “upgrade” weak majorization and containment from Jack to Macdonald polynomials.
Conjecture 7 (KT Conjecture for Macdonald polynomials).
The following are equivalent:
  1. (1)
    We have
    $\frac{P_{\lambda}(x+\bm{1};q,t)}{P_{\lambda}(\bm{1};q,t)}-\frac{P_{\mu}(x+\bm{1};q,t)}{P_{\mu}(\bm{1};q,t)}\in\mathbb{F}_{\geqslant 0}^{\mathbb{R}},\quad\forall x\in[0,\infty)^{n}.$ (5.11)
  2. (2)
    For some fixed $q_{0},t_{0}\in(1,\infty)$, we have
    $\frac{P_{\lambda}(x+\bm{1};q_{0},t_{0})}{P_{\lambda}(\bm{1};q_{0},t_{0})}-\frac{P_{\mu}(x+\bm{1};q_{0},t_{0})}{P_{\mu}(\bm{1};q_{0},t_{0})}\geqslant 0,\quad\forall x\in[0,\infty)^{n}.$ (5.12)
  3. (3)
    $\lambda$ weakly majorizes $\mu$.
As above, we now provide some positive evidence for this conjecture.
Theorem 5.8.
Each assertion in Conjecture 7 implies the next. Moreover, Conjecture 7 holds for all $n,\lambda$ when $\mu=(1^{m})$ for some $m\geqslant 1$.
Proof.
That $(1)\implies(2)$ is immediate, and that $(2)\implies(3)$ again follows via the argument used to show Theorem 3.2. Next, let $\lambda\succcurlyeq_{w}\mu$, and let $\mu^{\prime}={\bf 1}_{|\lambda|}$; then $\lambda\succcurlyeq\mu^{\prime}\supseteq\mu$. Now directly compute (suppressing $q_{0},t_{0}$ from the arguments):
$\frac{P_{\lambda}(x+{\bf 1})}{P_{\lambda}({\bf 1})}-\frac{P_{\mu}(x+{\bf 1})}{P_{\mu}({\bf 1})}=\left(\frac{P_{\lambda}(x+{\bf 1})}{P_{\lambda}({\bf 1})}-\frac{P_{\mu^{\prime}}(x+{\bf 1})}{P_{\mu^{\prime}}({\bf 1})}\right)+\left(\frac{P_{\mu^{\prime}}(x+{\bf 1})}{P_{\mu^{\prime}}({\bf 1})}-\frac{P_{\mu}(x+{\bf 1})}{P_{\mu}({\bf 1})}\right).$
At each $x\in[0,\infty)^{n}$, the first difference on the right is in $\mathbb{F}_{\geqslant 0}$ by Theorem 5.6, while the second is the scalar $M_{\mu^{\prime}}(x+{\bf 1})-M_{\mu}(x+{\bf 1})$, where both terms are positive reals. Now compute their ratio:
$\frac{M_{\mu^{\prime}}(x+{\bf 1})}{M_{\mu}(x+{\bf 1})}=\prod_{i=1}^{n}(x_{i}+1)^{\mu^{\prime}_{i}-\mu_{i}},$
and this is at least $1$ since all $x_{i}\in[0,\infty)$ and $\mu^{\prime}\supseteq\mu$. ∎
Conjecture 7 is not contingent upon our next and final conjecture, but the proof of its counterpart for Jack polynomials used a characterization of containment via Jack polynomials. Thus, we end with:
Conjecture 8 (Containment via Macdonald differences).
The Macdonald-counterpart over $\mathbb{F}_{\geqslant 0}$ of Theorem 1.1 holds.

### v2

@@ 3. $(2^{\prime})$
    For every $q_{0},t_{0}\in(0,1)$, we have
@@
    For some $q_{0},t_{0}\in(0,1)$, we have
@@
By the well-known duality $P_{\lambda}(x;q,t)=P_{\lambda}(x;1/q,1/t)$, one can define instead $\mathbb{K}_{\geqslant 0}\coloneqq\{\,f\in\mathbb{Q}(q,t)\mathclose{}\nonscript\;|\penalty 500\nonscript\;\mathopen{}f(q,t)\geqslant 0,\text{\ if\ }q,t\in(1,\infty)\,\}$ and get an equivalent statement.
@@
In Conjecture 5, it is clear that $(1)\Rightarrow(2)\&(2^{\prime})$, $(2)\Rightarrow(3)$, $(2^{\prime})\Rightarrow(3)$, and $(3)\Rightarrow(4)\Rightarrow(5)$. Hence the conjecture reduces to the implication $(5)\Rightarrow(1)$. Moreover, this implication holds for $\mu=(1^{|\lambda|})$; and it also holds for arbitrary $\lambda$ and $\mu$ in the case of two variables.
@@
Note that $(1)\Rightarrow(2)$ and $(2^{\prime})\Rightarrow(3)$ follows from Muirhead’s inequality; $(1)\Rightarrow(2^{\prime})$ and $(2)\Rightarrow(3)$ follow by specializing $q,t$; $(3)\Rightarrow(4)$ is obvious; $(4)\Rightarrow(5)$ is similar to Theorem 3.4 and is omitted.
Next, we assume $\mu=(1^{|\lambda|})$, with $n,\lambda$ arbitrary and show $(5)\Rightarrow(1)$. Haglund–Haiman–Loehr have shown [6] that the integral Macdonald polynomials are monomial-positive over $\mathbb{K}_{\geqslant 0}$ for $q,t\in(0,1)$, whence so are the $P_{\lambda}$. Now the calculation in Eq. 4.4 implies (1).
Finally, let $n=2$ and $\lambda,\mu$ have length at most 2. To show $(5)\Rightarrow(1)$, it suffices via telescoping to do so for majorization-adjacent partitions $\lambda=(d-i,i)$ and $\mu=(d-i-1,i+1)$. We showed this in Section 5 for $i=0$; and we use this to compute for $i>0$:
@@
$\in(x_{1}x_{2})^{i}\cdot\sum_{\nu\mathrel{\succcurlyeq}\xi}\mathbb{K}_{\geqslant 0}\cdot(M_{\nu}(x)-M_{\xi}(x))$
$=\sum_{\nu+(i,i)\mathrel{\succcurlyeq}\xi+(i,i)}\mathbb{K}_{\geqslant 0}\cdot(M_{\nu+(i,i)}(x)-M_{\xi+(i,i)}(x))\subseteq\mathcal{M}_{C}(\mathbb{K}_{\geqslant 0}).\qed$